OpenStudy (anonymous):

Which value of x makes the open sentence true? 5x – 6 = x + 12

OpenStudy (haseeb96):

5x-x= 12+6 4x=18 x=18\4 x=9\2 or x=4.5

OpenStudy (haseeb96):

let we prove it 5(9\2)-6=9\2+12 45-12\2=9+24\2 33\2=33\2 HEnce equation is being shows that x is 9\2